# EMMA: Scaling Mobile Manipulation via Egocentric Human Data

### v1

@@ IV-B Human and Robot Data Co-training
The goal of our system is to transfer knowledge from egocentric human mobile manipulation data ($D_{H}$) to scale policy performance, while still leveraging limited static robot manipulation data ($D_{R}$) for precise manipulation steps. The re-targeting and action alignment steps detailed in Sec. IV-A allow us to treat human and robot data as equal parts in a continuous spectrum of embodied data sources. But large domain gaps exist in the two data sources, in both sensing modalities and distributions. Thus, effectively learning from these data necessitates a unified learning framework capable of processing heterogeneous data sources. Inspired by recent works in cross-embodiment policy learning [46, 47, 48], we design an architecture based on a decoder-only Transformer with modality-specific input stems, a shared trunk, and multiple action and auxiliary output heads (Fig. 2).
@@ IV-C Auxiliary Phase Identification and Control Modulation
Phase Detection. We identify phases based on motion dynamics. For each frame, we compute the ratio of hand velocity to head velocity. Manipulation phases exhibit high hand-to-head velocity ratios ($\tau_{ratio}$) with low absolute head velocity ($\tau_{head}$), indicating dexterous hand motion while stationary. We filter candidate segments shorter than $\tau_{duration}$ frames to remove transient gestures.
@@
To spatially localize $K$ manipulation zones, we fit a Gaussian Mixture Model (GMM) with $K$ components to head positions during these high-ratio periods. Each frame is classified as manipulation (phase 0) if its probability density under the GMM exceeds $\tau_{pdf}$, otherwise navigation (phase 1). GMM provides direct probability density estimates needed for phase classification without requiring labeled training data, while being computationally efficient and interpretable for spatial clustering.
@@ V Experiments
Figure 4: Cumulative success rates across subtasks for three mobile manipulation tasks. EMMA (blue), trained without mobile teleoperation data, significantly outperforms Mobile ALOHA (orange) on Grocery Shopping and Handover Wine tasks (p << 0.05). Table Service variants show comparable performance. Error bars represent 95% confidence intervals from 50 trials.
@@
Handover Wine. The robot picks up a wine glass randomly placed within a 30cm×45cm area on a table. It turns right and navigates toward a human standing in a 3m×3m area. At a safe range, the robot hands the wine glass to the human’s right hand, testing precise navigation and state coverage transfer from human data.
@@
Baselines. We implement Mobile ALOHA [1] as our primary baseline to compare against teleoperated mobile robot data. Critically, it exemplifies the exact data collection paradigm—expensive mobile teleoperation—that EMMA aims to replace with human egocentric data. To ensure fair comparison, we modify it to the same HPT [47] backbone used in EMMA, isolating the comparison to data sources (human vs. teleoperation) rather than architectural differences. This baseline receives identical input modalities (egocentric RGB, wrist cameras, proprioception) and outputs robot joint actions $R_{q}\in\mathbb{R}^{14}$ plus base velocities $(v,\omega)\in\mathbb{R}^{2}$ recorded from the AgileX Tracer wheel encoders.
@@
Figure 5: For Handover Wine task, starting with a fixed amount of static manipulation data, we show that adding more human fullbody motion data for EMMA (blue) yields to greater performance gains compared to adding mobile robot teleoperation data collected under an equivalent amount of time for Mobile ALOHA (orange). The performance gap expands from 10% to 30% as data increases from 15 to 60 minutes.
Evaluation protocols. To ensure statistical significance of the results, We conduct 50 trials per task/model variant configuration, and record error bars with 95% confidence interval, while recording both success metrics and failure modes for qualitative analysis (Fig. 8). Each trial has a maximum duration of 2 minutes. For method comparison across all baseline and ablation experiments, we ensure identical environmental conditions and object positions. Alltogether, we conducted 950 mobile manipulation rollout evaluations to produce the results in this paper.
@@ V-A Main Results
EMMA achieves favorable performance compared to systems trained on teleoperated mobile robot data (H1).
Figure 6: EMMA co-trains static robot data in lab (left) with human demonstrations in a novel scene (right) to enable robot deployment in previously unseen environments.
@@
Figure 7: (a) EMMA generates to an unseen scene with 54% full task success rate. (b) Ablation study on the Handover Wine task. Removing either retargeting or phase switch causes significant performance drop.
Figure 8: Qualitative comparison of EMMA and Mobile ALOHA across three mobile manipulation tasks, showing representative success and failure modes.
@@
EMMA consistently outperforms teleoperation-based baselines across a range of mobile-manipulation tasks by leveraging egocentric human demonstrations (Fig. 4). In the Handover Wine task, with the same 100 demos of static robot data, replacing one hour of teleoperated mobile robot data with one hour of human mobile manipulation yields an 82% success rate—a 30% increase over teleoperated robot data, which had a 52% success rate. This improvement demonstrates the benefit of learning from egocentric human-human interaction. Leveraging egocentric human mobile manipulation data has also shown improvement in bimanual tasks found in the Grocery Shopping task. Lastly, using egocentric human mobile manipulation data has shown comparable performance to teleoperated mobile robot data on long-horizon Table Service task, a task that requires long-distance point-to-point navigation, where robot-teleoperated methods benefit from zero perspective gap. Taken together, these results show that egocentric human mobile manipulation data matches or even improves the value of equivalent robot teleoperation data and delivers safer, more reliable, and more scalable policies across diverse mobile-manipulation challenges.
@@
EMMA generalizes to unseen scenes. We evaluate scene generalization by testing EMMA in a novel environment seen only through human demonstrations. Specifically, we collect 30 minutes of human demonstrations in a new spatial layout where the wine recipient stands randomly within a larger 5m $\times$ 2m area (Fig. 6). Despite never seeing robot data from this environment, EMMA achieves 54% success rate (Fig. 7 (a)), demonstrating two key capabilities: (1) navigation behaviors learned from human data naturally adapt to the expanded spatial configuration, and (2) manipulation skills remain robust to visual domain shifts from the new environment. In contrast, Mobile ALOHA—trained exclusively on teleoperated data from the original environment is unable to complete the initial grasp due to the environmental changes (Fig. 8 (j)). This experiment validates that human data provides superior generalization through its inherent diversity compared to lab-constrained teleoperation.
@@
Ablation Studies (H2). Our ablation experiments on the Handover Wine task (Fig. 7 (b)) reveal the critical importance of both key components. Without kinematic retargeting, the success rate drops by 30%, as the robot frequently loses track of the human recipient when executing kinematically infeasible trajectories. The phase identification mechanism proves equally crucial—its removal causes complete task failure due to unintended base movements during grasping and unnecessary arm motions during navigation, resulting in collisions and dropped objects. These results validate that both human-robot embodiment alignment and learned phase modulation are essential for successful human-to-robot skill transfer in mobile manipulation.
@@
Human data scales mobile manipulation performance more efficiently compared to teleoperated robot data (H3). Keeping one hour of static robot manipulation data fixed, EMMA’s success rate climbs steadily from 0.36 to 0.82 as we increase human mobile manipulation data from 15 minutes to one hour. Mobile ALOHA also improves—from 0.26 to 0.52—when its robot teleoperation data grows from 15 minutes to one hour, but remains consistently below EMMA. This demonstrates that additional human mobile manipulation demonstrations yield substantially greater returns than equivalent increases in mobile robot teleoperation data. Qualitatively EMMA shows better robustness to target pose, fewer collisions with the environment and greater overall success. This is summarized in Fig. 8
VI Conclusion

### v2

@@ IV-B Human and Robot Data Co-training
Following the data processing and retargeting detailed in Sec. IV-A, which formally bridges the kinematic and distributional gaps, the goal of our system is to transfer knowledge from egocentric human mobile manipulation data ($D_{H}$) to scale policy performance, while still leveraging limited static robot manipulation data ($D_{R}$) for precise manipulation steps. But large domain gaps still exist in the two data sources, in both sensing modalities and distributions. Thus, effectively learning from these data necessitates a unified learning framework capable of processing heterogeneous data sources.
Inspired by recent works in cross-embodiment policy learning [liu2024rdt, wang2024hpt, ghosh2024octo], we design an architecture based on a decoder-only Transformer with modality-specific input stems, a shared trunk, and multiple action and auxiliary output heads (Fig. 2). Our goal is to transfer semantics of navigation and manipulation phases from human data. In addition, we want both the manipulation action prediction and navigation action prediction to benefit from large-scale human data. This motivates an architecture that shares a majority of the learnable parameters between the human and robot embodiments to promote shared representation, inspired by EgoMimic [egomimic].
@@ IV-C Auxiliary Phase Identification and Control Modulation
Phase Detection. We identify phases based on motion dynamics. For each frame, we compute the ratio of hand velocity to head velocity. Manipulation phases exhibit high hand-to-head velocity ratios ($\tau_{\text{ratio}}=2.0$) with low absolute head velocity ($\tau_{\text{head}}=0.4m/s$), indicating dexterous hand motion while stationary. We filter candidate segments shorter than $\tau_{\text{duration}}=30$ frames to remove transient gestures.
@@
To spatially localize $K$ manipulation zones, we fit a Gaussian Mixture Model (GMM) with $K=2$ components to head positions during these high-ratio periods. Each frame is classified as manipulation (phase 0) if its probability density under the GMM exceeds $\tau_{pdf}$, otherwise navigation (phase 1). GMM provides direct probability density estimates needed for phase classification without requiring labeled training data, while being computationally efficient and interpretable for spatial clustering. To assess accuracy, we hand-annotate 10% of uniformly sampled continuous segments from four tasks that require phase identification (see Section V). We report mean over frames (MoF), defined as the number of frames where the manual label matches the dataset label divided by the total number of sampled frames.
Task Table Service Handover Wine Handover Wine Scene Gen Grocery Shopping
MoF 0.962 0.965 0.929 0.963
@@ V Experiments
Figure 4: Cumulative success rates across subtasks for three mobile manipulation tasks. EMMA (blue), trained without mobile teleoperation data, significantly outperforms Mobile ALOHA (orange) on Grocery Shopping and Handover Wine tasks (p << 0.05). Table Service variants show comparable performance. Error bars represent 95% Clopper–Pearson confidence intervals with N = 50 trials.
@@
Handover Wine. The robot picks up a wine glass randomly placed on a tabletop dining mat. It turns right and navigates toward a human standing in a 3m×3m area. At a safe range, the robot hands the wine glass to the human’s right hand, testing precise navigation and state coverage transfer from human data.
@@
Push Chair. Grasp the backrest of a office chair and push it against a table. This task requires coupled arm-based motion to move and align the orientation of the chair. We establish a base EMMA model by co-training with 10 min of mobile robot data and 20 min of human demonstrations. We use this to: (1) Compare EMMA’s performance to Mobile ALOHA trained on 30 min (10+20) of pure robot data. (2) Compare our unified EMMA policy against an ablation using separate navigation/manipulation models.
Baselines. We implement Mobile ALOHA [fu2024mobile] as our primary baseline to compare against teleoperated mobile robot data. Critically, it exemplifies the exact data collection paradigm—expensive mobile teleoperation—that EMMA aims to replace with human egocentric data. To ensure fair comparison, we modify it to the same HPT [wang2024hpt] backbone used in EMMA, isolating the comparison to data sources (human vs. teleoperation) rather than architectural differences. This baseline receives identical input modalities (egocentric RGB, wrist cameras, proprioception) and outputs robot joint actions $R_{q}\in\mathbb{R}^{14}$ plus base velocities $(v,\omega)\in\mathbb{R}^{2}$ recorded from the AgileX Tracer wheel encoders.
@@
Figure 5: (a) For Handover Wine task, starting with a fixed amount of static manipulation data, we show that adding more human fullbody motion data for EMMA (blue) yields to greater performance gains compared to adding mobile robot teleoperation data collected under an equivalent amount of time for Mobile ALOHA (orange). The performance gap expands from 10% to 30% as data increases from 15 to 60 minutes. (b) EMMA generates to an unseen scene with 54% full task success rate.
Evaluation protocols. We conduct 50 trials per task/model variant, recording error bars with 95% Clopper-Pearson confidence interval with 50 trials. To compare the success rates between EMMA and our baseline, we use a two-proportion Z-test. We claim statistically significant outperformance for a model if p << 0.05. Each trial is limited to 2 minutes. Any collisions are recorded as a failure for the corresponding subtask phase. All experiments use identical environmental conditions and object positions. We conducted 1150 mobile manipulation rollout evaluations in total.
@@ V-A Main Results
Figure 6: Ablation study on the Handover Wine task. Removing either retargeting or phase switch causes significant performance drop.
Figure 7: Coupled arm-base motion task (Push Chair). Our unified EMMA (10m robot + 20m human) scales comparably to Mobile ALOHA (30m robot) and outperforms separate navigation/manipulation models, while excelling at the initial bimanual engagement.
Figure 8: Time-lapse of all five tasks (variants) showing the task flow. EMMA is capable of learning tasks from human-robot interaction to precise bimanual coordination.
Figure 9: Representative failure cases of Mobile ALOHA (a-c) and EMMA w/o retargeting (d). Analysis in Sec. V-A.
@@
EMMA achieves favorable performance compared to systems trained on teleoperated mobile robot data (H1). EMMA consistently outperforms teleoperation-based baselines across a range of mobile-manipulation tasks by leveraging egocentric human demonstrations (Fig. 4). This quantitative gap is reinforced by qualitative evidence, as the Mobile ALOHA model exhibits frequent critical failures (Figure 9): These include (a) navigation collisions, such as hitting the TV, and (b) flawed bimanual grasps, which result in incorrectly pushing the chair. In the Handover Wine task, with the same 50 demos of static robot data, replacing one hour of teleoperated mobile robot data with one hour of human mobile manipulation yields an 82% success rate—a 30% increase over teleoperated robot data, which had a 52% success rate. This improvement demonstrates the benefit of learning from egocentric human-human interaction. Similarly, for the Push Chair task that requires coupled arm-base movements, adding 20 min of human data to 10 min of mobile robot data allows EMMA to match the performance of Mobile ALOHA trained on 30 min of pure robot teleoperation, (Figure 7). Leveraging egocentric human mobile manipulation data has also shown improvement in bimanual tasks found in the Grocery Shopping task. Lastly, using egocentric human mobile manipulation data has shown comparable performance to teleoperated mobile robot data on long-horizon Table Service task, a task that requires long-distance point-to-point navigation, where robot-teleoperated methods benefit from zero perspective gap. Taken together, these results show that egocentric human mobile manipulation data matches or even improves the value of equivalent robot teleoperation data and delivers safer, more reliable, and more scalable policies across diverse mobile-manipulation challenges.
@@
EMMA generalizes to unseen scenes. We evaluate scene generalization by testing EMMA in a novel environment seen only through human demonstrations. Specifically, we collect 30 minutes of human demonstrations in a new spatial layout where the wine recipient stands randomly within a larger 5m $\times$ 2m area. Despite never seeing robot data from this environment, EMMA achieves 54% success rate (Fig. 5 (b)), demonstrating two key capabilities: (1) navigation behaviors learned from human data naturally adapt to the expanded spatial configuration, and (2) manipulation skills remain robust to visual domain shifts from the new environment. In contrast, Mobile ALOHA—trained exclusively on teleoperated data from the original environment is unable to complete the initial grasp due to the environmental changes (Fig. 5 (b)). This experiment validates that human data provides superior generalization through its inherent diversity compared to lab-constrained teleoperation.
@@
Ablation Studies (H2). Our ablation experiments on the Handover Wine task (Fig. 6 (a)) reveal the critical importance of both key components. Without kinematic retargeting, the success rate drops by 30%, as the robot frequently loses track of the human recipient when executing kinematically infeasible trajectories. The phase identification mechanism proves equally crucial—its removal causes complete task failure due to unintended base movements during grasping and unnecessary arm motions during navigation, resulting in collisions and dropped objects. These results validate that both human-robot embodiment alignment and learned phase modulation are essential for successful human-to-robot skill transfer in mobile manipulation. On the Push Chair task, a unified EMMA policy (10 min robot + 20 min human) outperforms separate navigation and manipulation models on precise bimanual chair grasp and continuous engagement. (Figure 7), indicating that single-policy co-training better exploits human data for arm-base coupling. These results validate that our design choices are essential for successful human-to-robot skill transfer. Our observed representative failure modes further reinforce this (Figure 9): (c) without phase switching, Mobile ALOHA makes unintended base motions that hinder precise manipulation. (d) Without kinematic retargeting, executing raw human waypoints yields infeasible trajectories;
@@
Human data scales mobile manipulation performance more efficiently compared to teleoperated robot data (H3). Keeping one hour of static robot manipulation data fixed, EMMA’s success rate climbs steadily from 0.36 to 0.82 as we increase human mobile manipulation data from 15 minutes to one hour. Mobile ALOHA also improves—from 0.26 to 0.52—when its robot teleoperation data grows from 15 minutes to one hour, but remains consistently below EMMA. This demonstrates that additional human mobile manipulation demonstrations yield substantially greater returns than equivalent increases in mobile robot teleoperation data.
VI Limitation
Despite the strengths of EMMA, our approach inherits several inherent limitations from learning mobile manipulation primarily through human demonstrations. First, the framework assumes that the visual and spatial distributions encountered during robot deployment lie within, or close to, those seen in human demonstrations. This assumption may break down when the kinematic or viewpoint differences between human and robot embodiments become substantial. Second, the direct transfer of human navigation strategies is insufficient for tasks requiring tightly coupled arm–base coordination. Nevertheless, we showed that human data provided significant scaling when combined with even limited mobile robot data.
VII Conclusion
@@ VIII Acknowledgment
-A Problem Formulation Details
-A1 Bi-source Imitation Learning Problem
We formalize our mobile manipulation learning problem as imitation learning from two heterogeneous data sources: human mobile manipulation demonstrations $D_{H}$ and static robot manipulation demonstrations $D_{R}$. The observation spaces are:
  • $\mathcal{O}^{H}$: Human observation space containing egocentric RGB images $I_{\text{ego}}$, hand poses $H_{p}\in SE(3)\!\times\!SE(3)$ in the egocentric RGB camera frame, and 3D head pose $H_{d}\in SE(3)$.
  • $\mathcal{O}^{R}$: Robot observation space containing egocentric RGB $I_{\text{ego}}$, wrist camera images $I_{\text{wrist}}$, joint positions $R_{q}\in\mathbb{R}^{14}$, and end-effector poses $R_{p}\in SE(3)\!\times\!SE(3)$.
The shared egocentric view $I_{\text{ego}}$ serves as the primary linking modality between embodiments. During deployment, we assume:
  • Manipulation phase: $\mathcal{O}_{\text{deploy}}\subseteq\mathcal{O}^{R}\cup\mathcal{O}^{H}$
  • Navigation phase: $\mathcal{O}_{\text{deploy}}\subset\mathcal{O}^{H}$
This assumption enables direct transfer of navigation knowledge purely from human data while leveraging both sources for manipulation.
-B Phase-Conditioned Policy
The policy operates under implicit phase conditioning:
$\pi(a_{t}\mid o_{t},\phi_{t}),$ (6)
where $\phi_{t}\in\{0,1\}$ indicates manipulation (0) or navigation ($1$). Phase identification uses unsupervised clustering with parameters:
  • Velocity ratio threshold: $\tau_{\text{ratio}}=2.0$
  • Head velocity threshold: $\tau_{\text{head}}=0.4~\text{m/s}$
  • Minimum phase duration: $\tau_{\text{duration}}=30$ frames (1 s at 30 fps)
  • GMM components: $K=2$ for spatial manipulation-zone identification
  • Probability density threshold: $\tau_{\text{pdf}}=10^{-3}$
-C Differential-Drive Retargeting Constraints
The optimization problem (Equations 1–4 in the main text) employs the following bounds and weights:
  • Velocity constraints: $v\in[-1.6,\,1.6]~\text{m/s}$, $\omega\in[-1.5,\,1.5]~\text{rad/s}$
  • Optimization weights: $\lambda_{\text{pos}}=25$, $\lambda_{\text{yaw}}=2$, $\lambda_{\text{smooth}}=1$
  • Integration timestep: $\Delta t=0.02\times\text{STEP}$ seconds
  • Navigation sampling step: $\text{STEP}_{\text{nav}}=8$ frames in all trained models.
The retargeting ensures kinematic feasibility while minimizing deviation from human trajectories through constrained optimization.
-D Action Chunking and Temporal Alignment
Actions are sampled at different temporal resolutions and then unified:
Modality Horizon Sampling Step Interpolated Length
Navigation 10 8 frames 100
Manipulation 10 4 frames 100
We interpolate to a fixed chunk length of 100 to ensure consistent action representation across modalities. Navigation actions undergo phase-aware filtering in which manipulation-phase waypoints are replaced with interpolated trajectories to the first navigation waypoint.
-E Coordinate Frame Alignment
All action predictions are expressed in the egocentric camera frame at observation time:
  • Human hand poses: Transformed using the SLAM-estimated camera pose $T_{\text{ego}}^{-1}$.
  • Robot end-effector: Transformed using the calibrated hand–eye transformation.
  • Waypoint history: Maintained as $\tilde{W}_{t}=\{\,T_{\text{ego}}^{-1}\!\cdot\!h_{\text{base}}^{\,t-k_{i}}\,\}_{i=1}^{K_{h}}$ with displacement threshold $d_{\text{thresh}}=0.25$ m.
Z-score normalization is applied independently per data source ($D_{H}$, $D_{R}$) to handle distributional differences.
-F Data Quality Constraints
The framework enforces several quality constraints:
  • Hand-tracking confidence filtering: Frames with confidence <0<0 are excluded, allowing data collectors to temporarily lift the glasses to reset a scene without corrupting supervision.
  • Waypoint sparsity: New waypoints are added only when $\|h_{\text{base}}^{t}-h_{\text{base}}^{t-k}\|_{2}\geq d_{\text{thresh}}$, enforcing spatial sampling at $\geq d_{\text{thresh}}$ increments.
  • Phase-transition smoothing: Segments shorter than $\tau_{\text{duration}}$ (30 frames) are filtered.
-G Temporal Consistency Assumption
Phase transitions are assumed sparse, with manipulation zones spatially localized. This assumption—enforced via minimum-duration constraints and GMM-based spatial clustering—enables reliable phase prediction and prevents oscillatory switching during deployment.
-H Network Architecture
Architecture
Each input modality is encoded with a dedicated stem (e.g. a ResNet-18 for images and an MLP for low-dimensional state). Each stem produces a feature sequence that is tokenized via cross-attention using a small bank of learnable queries, yielding a fixed number of latent tokens per modality. Sinusoidal positional embeddings are added (i) within each modality stream prior to tokenization and (ii) again after concatenating all modality tokens. We then append a learnable bank of action tokens whose count matches the action horizon, and process the full sequence with a Transformer trunk (Fig. 10).
Figure 10: Detailed network architecture comparison of Mobile ALOHA (left) and EMMA (right). We supervise all heads with Smooth-$\ell_{1}$ ($\beta{=}0.05$) + MSE loss.
TABLE I: Hyperparameters for EMMA vs. Mobile ALOHA.
Item EMMA Mobile ALOHA
Input resolution $H\times W\times C$ $480\times 640\times 3$ $480\times 640\times 3$
Action chunk length $k$ $100$ $100$
Image encoder ResNet-18 (ImageNet init.) ResNet-18 (ImageNet init.)
Modality embedding dim $d_{\text{proj}}$ $256$ $256$
Latent queries per modality $L$ $16$ $16$
Stem cross-attn heads $D_{\text{stem}}$ 88 88
Stem per-head dim $d_{\text{attn}}$ $64$ $64$
Transformer trunk hidden dim $d$ $256$ $256$
Transformer trunk heads $D_{\text{trunk}}$ 88 88
Transformer trunk layers $N_{\text{trunk}}$ $16$ $16$
Learnable action queries $100$ $100$
Positional encoding Sinusoidal (per-stream & global) Sinusoidal (per-stream & global)
Prediction head MLP (hidden width $512$, SiLU, LayerNorm) MLP (hidden width $512$, SiLU, LayerNorm)
Supervision loss Smooth-$\ell_{1}$ ($\beta{=}0.05$) + MSE Smooth-$\ell_{1}$ ($\beta{=}0.05$) + MSE
Nav. upsampling $10\!\rightarrow\!100$ (linear $x,y$; $\operatorname{atan2}$ yaw)
Action dims $d_{a}$ 77 or $14$ (robot joints) + $6$ or $12$ (human ee pose) + 3 (nav waypoints) 77 or $14$ (robot joints) + $2$ (base velocities)
Training targets and loss
All targets (e.g. manipulation trajectories, navigation waypoints, phase) are supervised directly with a smooth-$\ell_{1}$ objective. When a shared head is active it is supervised with the corresponding shared target; auxiliary heads (when present) are supervised on their respective labels using the same backbone features.
-I Baseline Details
As discussed in Sec. V of the main text, we compare EMMA with Mobile ALOHA by varying the source of mobile-manipulation supervision. Like EMMA, Mobile ALOHA co-trains the same static robot manipulation data (teleoperation) with an additional source of mobile data (Fig. 10). However, Mobile ALOHA replaces human full-body data with mobile teleoperation, where the robot’s base linear and angular velocities are recorded at 50 Hz using wheel encoders on the AgileX Tracer platform. For a fair comparison, the stem and trunk architectures remain identical to EMMA, while the prediction head is extended to output a chunk of future base velocities (Table I).
To maximize data collection efficiency, two teleoperators are employed: one controls the robot arms using a leader–follower system physically connected to the robot laptop via USB extension cables, while the other simultaneously drives the mobile base using a remote controller.